# MoCA: Multi-modal Cross-masked Autoencoder for Digital Health Measurements

### v1

@@ 3.1 Tokenization
[…]
@@ 4 Theoretical Results
We present here theoretical results on the bounds of the MAE loss function under different random masking schemes. Additionally, we establish a connection between the MAE loss function bounds and canonical correlation analysis (CCA) using a Reproducing Kernel Hilbert Space (RKHS) approach.
@@ 4.1 Notation
Let $\bar{x}$ denote a generalized “image” from a dataset that consists of a matrix of data patches of multiple modalities. Each row of $\bar{x}$ corresponds to one modality. Given a total of $C$ modalities, suppose each modality contains $P$ patches, where each patch is a time series of a fixed length. Then, each image consists of a total of $CP$ patches.
[…]
Given a mask ratio $\rho$, a realization of the random masking matrix $\delta$ is a matrix taking values of 0 or 1, where the total number of ones is $\lfloor\rho CP\rfloor$. Conditional on $\bar{x}$, the probability distribution of $\delta$ over all admissible masking matrices, subject to the constraints of the sampling scheme, determines the distribution of masked and unmasked views.
4.2 Bounds for MAE Reconstruction Losses
The Masked Autoencoder (MAE) model consists of an encoder $f$, which maps the unmasked view $x^{U}$ to a latent feature $z^{U}=f(x^{U})$, and a decoder $g$, which maps $z^{U}$ back to the space of the masked view $x^{M}$. The objective of the MAE is to minimize the loss function
$L_{\text{MAE}}=\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\|g\circ f(x^{U})-x^{M}\|^{2}.$
We assume that $x^{M}$, $x^{U}$, and the reconstructed masked view $\hat{x}^{M}:=g\circ f(x^{U})$ are all $l_{2}$-normalized to the same scale.
@@ 4.2 Bounds for MAE Reconstruction Losses
For any non-degenerate decoder $g$, assume there exists a pseudo-inverse encoder $f_{g}$ such that the pseudo autoencoder $h_{g}:=g\circ f_{g}$ satisfies $\mathbb{E}_{x}\|h_{g}(x)-x\|^{2}\leq\epsilon$, where $x$ represents either unmasked view $x^{U}$ or masked view $x^{M}$ and $\epsilon$ is a small positive constant.
It has been shown in (Zhang et al., 2022b) that given Assumption 4.1, the MAE loss is bounded from below:
@@
$L_{\text{MAE}}(h)\geq-\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\langle h(x^{U}),h_{g}(x^{M})\rangle-\epsilon+\text{const},$ (1)
[…]
4.3 Kernel Representations
We represent the encoder and decoder mappings using Reproducing Kernel Hilbert Space (RKHS) representations. Assume there exist mappings $\Phi$ and $\tilde{\Phi}$ such that
$h(x^{U})=g\circ f(x^{U})=[\Phi(x_{c_{i_{1}},p_{i_{1}}}),\dots,\Phi(x_{c_{i_{n_{u}}},p_{i_{n_{u}}}})],$ (2)
where $\{i_{1},\dots,i_{n_{u}}\}$ are indices of the unmasked patches, and $n_{u}=CP-\lfloor\rho CP\rfloor$ is the total number of patches in the unmasked view $x^{U}$, and similarly,
$h_{g}(x^{M})=g\circ f_{g}(x^{M})=[\tilde{\Phi}(x_{c_{j_{1}},p_{j_{1}}}),\dots,\tilde{\Phi}(x_{c_{j_{n_{m}}},p_{j_{n_{m}}}})],$ (3)
where $\{j_{1},\dots,j_{n_{m}}\}$ are indices of the masked patches, and $n_{m}=\lfloor\rho CP\rfloor$ is the total number of patches in the masked view $x^{M}$.
Using the kernel trick, we can further write (2) using the representation
$\Phi(x)=\sum_{k}\alpha_{u}^{T}(k)k_{u}(x,x_{c_{i_{k}},p_{i_{k}}}),$
where $k_{u}(\cdot,\cdot)$ is the kernel associated with the RKHS in which $\Phi$ resides in, so that for all pairs of unmasked patches $(x_{c_{i_{k}},p_{i_{k}}},x_{c_{i_{l}},p_{i_{l}}})$,
$k_{u}(x_{c_{i_{k}},p_{i_{k}}},x_{c_{i_{l}},p_{i_{l}}})=\langle\Phi(x_{c_{i_{k}},p_{i_{k}}}),\Phi(x_{c_{i_{l}},p_{i_{l}}})\rangle,$
and $\alpha_{u}$ is a vector ($\alpha^{T}_{u}$ is its transpose) in which the $k$th element $\alpha_{u}(k)$ is the weight of the basis spanned by the reproducing kernel at unmasked patch $x_{c_{i_{k}},p_{i_{k}}}$.
Similarly,
$\tilde{\Phi}(x)=\sum_{l}\alpha_{m}^{T}(l)k_{m}(x,x_{c_{j_{l}},p_{j_{l}}}),$ (4)
where $k_{m}(\cdot,\cdot)$ is the kernel associated with the RKHS in which $\tilde{\Phi}$ resides in, so that for all pairs of masked patches, $(x_{c_{j_{k}},p_{j_{k}}},x_{c_{j_{l}},p_{j_{l}}})$,
$k_{m}(x_{c_{j_{k}},p_{j_{k}}},x_{c_{j_{l}},p_{j_{l}}})=\langle\tilde{\Phi}(x_{c_{j_{k}},p_{j_{k}}}),\tilde{\Phi}(x_{c_{j_{l}},p_{j_{l}}})\rangle,$
and $\alpha_{m}$ is a vector, in which the $l$th element $\alpha_{m}(l)$ is the weights of the basis spanned by the reproducing kernel at masked patch $x_{c_{j_{l}},p_{j_{l}}}$.
4.4 Connection to Canonical Correlation Analysis (CCA)
Minimizing the lower bound in (1) is equivalent to maximizing the quantity $\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\langle h(x^{U}),h_{g}(x^{M})\rangle$, which, using the RKHS representations in the previous section, can be written as
$\begin{split}\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}\langle[\Phi(x_{c_{i_{1}},p_{i_{1}}}),\dots,\Phi(x_{c_{i_{n_{u}}},p_{i_{n_{u}}}})],\\ [\tilde{\Phi}(x_{c_{j_{1}},p_{j_{1}}}),\dots,\tilde{\Phi}(x_{c_{j_{n_{m}}},p_{j_{n_{m}}}})]\rangle,\end{split}$
which equals to
$\mathbb{E}_{\bar{x}}\mathbb{E}_{x^{U},x^{M}\mid\bar{x}}[\alpha_{u}^{T}K_{u}K_{m}\alpha_{m}],$ (5)
where $K_{u}=k_{u}-k_{u}\mathbf{1}-\mathbf{1}k_{u}+\mathbf{1}k_{u}\mathbf{1}$ with $k_{u}$ denoting the matrix whose $(k,l)$th element is $k_{u}(x_{c_{i_{k}},p_{i_{k}}},x_{c_{i_{l}},p_{i_{l}}})$ and $\mathbf{1}$ denotes the matrix with all elements equal to one. The definition of $K_{m}$ follows the same pattern.
The representation of (5) has a close connection to the Canonical Correlation Analysis (CCA). Given a distribution of random images ${\bar{x}}$, the random masking scheme, characterized by the distribution of $\delta$, determines the distribution of unmasked and masked view assignments. For each pair $(x^{U},x^{M})$, the matrices $k_{u}$ and $k_{m}$ are determined by their respective compositions $x^{U}=(x_{c_{i_{1}},p_{i_{1}}},\dots,x_{c_{i_{n_{u}}},p_{i_{n_{u}}}})$ and $x^{M}=(x_{c_{j_{1}},p_{j_{1}}},\dots,x_{c_{j_{n_{m}}},p_{j_{n_{m}}}})$.
Given a realization of $\delta$, the resulting $(x^{U},x^{M})$, and under the assumption of $l_{2}$ regularization of both unmasked and masked views, the optimal solution to (5) is obtained by finding vectors $\alpha^{*}_{u}$ and $\alpha^{*}_{m}$ that satisfy
$(\alpha^{*}_{u},\alpha^{*}_{m})=\operatorname*{arg\,max}_{\alpha_{u},\alpha_{m}}\frac{\alpha^{T}_{u}K_{u}K_{m}\alpha_{m}}{\sqrt{(\alpha^{T}_{u}K_{u}^{2}\alpha_{u})(\alpha^{T}_{m}K_{m}^{2}\alpha_{m})}},$ (6)
which can be solved using the Lagrangian multiplier (for details, see, for example, (Yang et al., 2019; Andrew et al., 2013)).
A special case is when the mappings $\Phi$ and $\hat{\Phi}$ are linear, then the objective in (6) becomes the classical CCA objective
$\max_{\alpha_{u},\alpha_{m}}\frac{\alpha^{T}_{u}\Sigma_{UM}\alpha_{m}}{\sqrt{(\alpha^{T}_{u}\Sigma_{UU}\alpha_{u})(\alpha^{T}_{m}\Sigma_{MM}\alpha_{m})}},$ (7)
where $\Sigma_{UM}$ denotes the cross covariance matrix between all pairs of unmasked and masked patches, $\Sigma_{UU}$ denotes the cross covariance matrix of all pairs of unmasked patches, and $\Sigma_{MM}$ denotes the cross covariance matrix of all pairs of masked patches. The solution to this problem will be discussed in the next section.

### v2

@@ 3.1 Tokenization
Consider a multi-modal time series sample $x\in\mathbb{R}^{1\times L\times C}$, where $C$ represents the number of modalities and $L$ denotes the sequence length of each modality. Since the computational complexity of Transformer-based models scales with sequence length $L$, processing raw time series data without tokenization becomes computationally prohibitive. To address this, we segment each modality into $P=\lfloor L/L_{p}\rfloor$ ( $\lfloor\cdot\rfloor$ denotes the floor function) non-overlapping patches (time windows), where $L_{p}$ defines the patch size (window length). We adopt the channel-independent tokenization, building upon the framework introduced by (Nie et al., 2023):
@@ 4 Theoretical Results
In this section, we outline a theoretical analysis of the proposed method and establish a connection between the MAE reconstruction loss during pretraining and Kernel Canonical Correlation Analysis (KCCA) within a Reproducing Kernel Hilbert Space (RKHS) framework.
@@ 4.1 Notation
Let $\{x_{i}:i=1,\dots,n\}$ denote the set of $n$ generalized ”images” from a dataset. In our setting, an image is a 4-second window of multi-modal time-series data recorded by one or multiple concurrently worn wearable devices. Suppose there are $C$ modalities (rows), each split into $P$ contiguous, non-overlapping patches of fixed length $L_{p}$ (columns). Hence every image contains $C\times P$ patches. We write the $(c,p)$-th patch as $x_{i,(c,p)}\in\mathbb{R}^{L_{p}}$, where $c=1,\dots,C$ and $p=1,\dots,P$.
For the $i$th image, let $m_{i}\in\{0,1\}^{C\times P}$ be the binary masking matrix, with entries $m_{i,(c,p)}$. A value of $1$ indicates that all elements in the corresponding patch are masked and 0 indicates elements unmasked. The masked and unmasked “views” of $x_{i}$ are then defined as collections of all masked and unmasked patches in the image, respectively:
$x_{i}^{M}=x_{i}\odot m_{i},\qquad x_{i}^{U}=x_{i}\odot(\mathbf{1}-m_{i}),$
where $\odot$ denotes patch-wise (Hadamard) multiplication and $\mathbf{1}$ is the all-ones matrix of the same shape as $m_{i}$.
Fix a mask ratio $\rho\in(0,1)$. For all $i$, realizations of $m_{i}$ must satisfy the constraint that
$\sum_{c=1}^{C}\sum_{p=1}^{P}m_{i,(c,p)}=\lfloor\rho\,CP\rfloor.$
Let
$\mathcal{M}_{\rho}=\Bigl{\{}m\in\{0,1\}^{C\times P}:\textstyle\sum_{c,p}m_{(c,p)}=\lfloor\rho\,CP\rfloor\Bigr{\}}$
be the set of admissible masks under mask ratio $\rho$. Conditional on $x_{i}$, the sampling scheme specifies a probability measure $P(\cdot\mid x_{i})$ supported on $\mathcal{M}_{\rho}$; drawing $m_{i}\sim P(\cdot\mid x_{i})$ independently for each $i$ determines which patches form the masked view $x_{i}^{M}$ and which remain in the unmasked view $x_{i}^{U}$. Additional structural constraints can be imposed by further restricting the support of $P$. For example, in a synchronized masking scheme, all modalities sharing the same time interval receive the same mask indicator, so an entire temporal slice is either fully masked or fully unmasked across modalities.
4.2 Bound for MAE Reconstruction Loss
The Masked Autoencoder (MAE) model consists of an encoder $f$ and a decoder $g$. Given a masking matrix $m_{i}$ and its induced partition $(x_{i}^{U},x_{i}^{M})$, the encoder $f$ maps the unmasked view to a latent feature $z_{i}^{U}$, which the decoder $g$ then maps to a reconstructed masked view:
$z_{i}^{U}=f\!\left(x_{i}^{U}\right),\qquad\hat{x}_{i}^{M}=g\!\left(z_{i}^{U}\right)=(g\!\circ\!f)\!\left(x_{i}^{U}\right),$
where we define $h:=g\!\circ\!f$ for convenience.
Assumption 4.1 (normalization).
All views and reconstructions are $\ell_{2}$-normalized:
$\|x_{i}^{U}\|_{2}=\|x_{i}^{M}\|_{2}=\|\hat{x}_{i}^{M}\|_{2}=1,\quad\text{for all }i.$
The loss of the MAE is the expected mean-squared reconstruction error on the masked views:
$L_{\text{MAE}}=\mathbb{E}_{x_{i}}\mathbb{E}_{m_{i}\sim P(\cdot\mid x_{i})}\Bigl{[}\bigl{\|}\bigl{(}g\!\circ\!f\bigr{)}(x_{i}^{U})-x_{i}^{M}\bigr{\|}_{2}^{2}\Bigr{]},$ (1)
where the inner expectation is taken with respect to the masking distribution $P$ introduced earlier. In practice (1) is approximated by a Monte Carlo average over minibatches of samples and independently drawn masks. Hence, we write the empirical distribution version of the MAE loss function as:
$\widehat{L}_{\text{MAE}}=\frac{1}{n}\sum_{i=1}^{n}\bigl{\|}\bigl{(}g\!\circ\!f\bigr{)}\!\bigl{(}x_{i}^{U}\bigr{)}-x_{i}^{M}\bigr{\|}_{2}^{2}.$ (2)
@@ 4.2 Bound for MAE Reconstruction Loss
For any non-degenerate decoder $g$, there exists a pseudo-inverse encoder $f_{g}$ such that the pseudo autoencoder $h_{g}:=g\circ f_{g}$ satisfies $\mathbb{E}_{x}\|h_{g}(x)-x\|_{2}^{2}\leq\epsilon$, where $x$ denotes either the unmasked view $x^{U}$ or the masked view $x^{M}$ and $\epsilon>0$ is a small constant.
Remark 4.3.
Assumption 4.2 requires that, for any fixed decoder $g$, there exists a pseudo-inverse encoder $f_{g}$ that produces latent features which $g$ can map back to a close approximation of the original input $x$ (masked or unmasked view). The reconstruction need not be exact—only within a small mean-squared error $\epsilon$.
It was shown in (Zhang et al., 2022b) that under Assumption 4.2, the MAE loss has a lower bound:
@@
$L_{\mathrm{MAE}}\;\geq\;-\mathbb{E}_{x}\,\mathbb{E}_{m\sim P(\cdot\mid x)}\bigl{\langle}h(x^{U}),h_{g}(x^{M})\bigr{\rangle}-\varepsilon+\text{const},$ (3)
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product. The constant term depends only on the data distribution and the masking scheme, not on the learnable encoder $f$ and decoder $g$. Therefore, minimizing this bound is equivalent to maximizing the alignment term $\mathbb{E}_{x}\mathbb{E}_{m\sim P(\cdot\mid x)}\langle h(x^{U}),h_{g}(x^{M})\rangle$. Given $n$ samples and independently drawn masks $\{m_{i}\}$, the alignment term is approximated by the empirical average
$\frac{1}{n}\sum_{i=1}^{n}\langle h(x_{i}^{U}),h_{g}(x_{i}^{M})\rangle.$ (4)
We next adopt an RKHS representation of the encoder–decoder mappings to express the maximization of this empirical alignment term as an equivalent kernel canonical correlation analysis (KCCA) objective.
4.3 Reproducing Kernel Hilbert Space Representations
Let $k_{u}$ be a positive-definite kernel that measures the similarity between any two unmasked views in $\{x^{U}_{i}\}$. Similarly, let $k_{m}$ be a positive-definite kernel that measures the similarity between any two masked views in $\{x^{M}_{i}\}$. Note that here both $k_{u}$ and $k_{m}$ are defined in the generalized case of kernels for vector-valued functions. That is, $k_{u}$ and $k_{m}$ are matrix-valued (operator-valued) positive-definite kernels: for any pair of inputs they return a $d\times d$ positive-definite matrix, where $d\geq 1$ is the dimensionality of the vectors $h(x_{i}^{U})$ and $h_{g}(x_{i}^{M})$.
These kernels induce vector-valued Reproducing Kernel Hilbert Spaces (RKHS) $\mathcal{H}_{U}$ and $\mathcal{H}_{M}$, whose elements are functions defined on the unmasked and masked views, respectively. In certain settings one may adopt a shared similarity measure by taking $k_{u}=k_{m}$. For generality we allow here the two views to be equipped with distinct kernels.
Remark 4.4.
When each view is composed of multiple patches, the view-level kernels $k_{u}$ and $k_{m}$ can be constructed from patch-level kernels via a mean embedding. Let $\mathcal{P}(x_{i}^{U})$ denote the set of patches in the unmasked view $x_{i}^{U}$, and let $k_{\mathrm{patch}}$ be a positive-definite kernel between individual patches. The corresponding view-level kernel can be defined as
$k_{u}(x_{i}^{U},x_{j}^{U})=\frac{1}{|\mathcal{P}(x_{i}^{U})|\,|\mathcal{P}(x_{j}^{U})|}\sum_{p\in\mathcal{P}(x_{i}^{U})}\sum_{q\in\mathcal{P}(x_{j}^{U})}k_{\mathrm{patch}}(p,q),$
i.e., the average of all patch–patch kernel evaluations between the two views. An analogous definition applies to $k_{m}$ for masked views. Equivalently, if $\psi_{\mathrm{patch}}(\cdot)$ is the feature map associated with $k_{\mathrm{patch}}$, the view-level feature embedding can be taken as the mean patch embedding
$\psi_{\mathrm{view}}(x_{i}^{U})=\frac{1}{|\mathcal{P}(x_{i}^{U})|}\sum_{p\in\mathcal{P}(x_{i}^{U})}\psi_{\mathrm{patch}}(p),$
so that $k_{u}(x_{i}^{U},x_{j}^{U})=\langle\psi_{\mathrm{view}}(x_{i}^{U}),\,\psi_{\mathrm{view}}(x_{j}^{U})\rangle$. This mean-embedding construction preserves positive-definiteness and naturally aggregates patch-level similarities into a view-level similarity.
By the Representer Theorem (Schölkopf et al., 2001; Micchelli & Pontil, 2005; Álvarez et al., 2012), given Assumption 4.1, any $h\in\mathcal{H}_{U}$ and $h_{g}\in\mathcal{H}_{M}$ that maximize (4) have the following representations:
$h(x_{i}^{U})$ $=\sum_{k=1}^{n}k_{u}(x_{i}^{U},x_{k}^{U})\alpha_{k},$ (5)
$h_{g}(x_{i}^{M})$ $=\sum_{k=1}^{n}k_{m}(x_{i}^{M},x_{k}^{M})\beta_{k},$
with coefficient vectors $\alpha_{k},\beta_{k}\in\mathbb{R}^{d}$ for $k=1,\dots,n$. When $d=1$ (scalar outputs), these reduce to the familiar scalar-RKHS representations: weighted linear combinations of $\{k_{u}(\,\cdot\,,x_{k}^{U})\}$ and $\{k_{m}(\,\cdot\,,x_{k}^{M})\}$.
Remark 4.5.
In this formulation we assume that the learnable mappings $h=g\!\circ\!f$ and $h_{g}=g\!\circ\!f_{g}$ can be well-approximated by elements of $\mathcal{H}_{U}$ and $\mathcal{H}_{M}$, respectively. In practice, this requires choosing kernels $k_{u}$ and $k_{m}$ whose associated RKHS are rich enough to capture the views’ latent patterns as well as the similarities in semantics among the views, so that the encoder–decoder mappings are representable in the RKHS form (5). For kernel choice, see operator-valued kernels for vector-valued functions and multi-task learning (Micchelli & Pontil, 2005; Carmeli et al., 2010; Álvarez et al., 2012), and alignment/deformation-tolerant time series kernels such as (Cuturi et al., 2007; Anden & Mallat, 2014; Salvi et al., 2021).
4.4 Connection to Kernel Canonical Correlation Analysis
Let $K_{U}\in\mathbb{R}^{nd\times nd}$ denote the block Gram matrix whose $(i,j)$-th block $K_{U}^{(i,j)}\in\mathbb{R}^{d\times d}$ is given by $k_{u}(x_{i}^{U},x_{j}^{U})$. Similarly, let $K_{M}\in\mathbb{R}^{nd\times nd}$ denote the block Gram matrix whose $(i,j)$-th block is $k_{m}(x_{i}^{M},x_{j}^{M})$. When $d=1$, $K_{U}$ and $K_{M}$ reduce to the standard $n\times n$ scalar-valued Gram matrices.
With the representations (5), the empirical alignment term (4) can be expressed as
$\frac{1}{n}\,\alpha^{\!\top}K_{U}K_{M}\beta,$ (6)
where $\alpha=[\alpha_{1},\dots,\alpha_{n}]^{\!\top}$ and $\beta=[\beta_{1},\dots,\beta_{n}]^{\!\top}$. Under Assumption 4.1 and omitting the constant factor $1/n$, the optimal solution to (6) is obtained by solving
$(\alpha^{*},\beta^{*})=\arg\max_{\alpha,\beta}\frac{\alpha^{\!\top}K_{U}K_{M}\beta}{\sqrt{\alpha^{\!\top}K_{U}^{2}\alpha\;\beta^{\!\top}K_{M}^{2}\beta}},$ (7)
which is the uncentered kernel canonical correlation analysis (KCCA) objective. When centering is appropriate (as in the example in the next section), one can replace each Gram matrix by its centered form $\tilde{K}=HKH$ with $H=I-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\top}$, which corresponds to centering in feature space and yields the classical centered KCCA formulation. For more details, see, e.g., (Schölkopf et al., 2001; Fukumizu et al., 2007; Andrew et al., 2013).
The KCCA problem is solved using Lagrange multipliers. With the normalization constraints $\alpha^{\top}K_{U}^{2}\alpha=1$ and $\beta^{\top}K_{M}^{2}\beta=1$, the stationary equations from (7) can be written as
$K_{U}K_{M}\beta=\rho\,K_{U}^{2}\alpha,\qquad K_{M}K_{U}\alpha=\rho\,K_{M}^{2}\beta.$
These can be combined into the block generalized eigenvalue problem
$\begin{bmatrix}0&K_{U}K_{M}\\ K_{M}K_{U}&0\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}=\rho\begin{bmatrix}K_{U}^{2}&0\\ 0&K_{M}^{2}\end{bmatrix}\begin{bmatrix}\alpha\\ \beta\end{bmatrix}$. (8)
In practice, regularization is applied to avoid degeneracy by replacing $K_{U}^{2}$ and $K_{M}^{2}$ in (8) with
$K_{U}^{2}+\gamma_{U}K_{U},\qquad K_{M}^{2}+\gamma_{M}K_{M},$
where $\gamma_{U},\gamma_{M}>0$ are small regularization parameters. The leading eigenvalue $\rho$ gives the maximum objective, and the associated eigenvectors yield the optimal coefficients $\alpha^{*}$ and $\beta^{*}$. For detailed derivations, see, e.g., (Hardoon et al., 2004; Fukumizu et al., 2007; Andrew et al., 2013; Yang et al., 2019).
4.5 Influence of Masking Schemes
A stochastic masking policy is specified by a probability measure $P(\,\cdot\,|x)$ on the admissible set $\mathcal{M}_{\rho}$. For each image $x_{i}$, a mask $m_{i}\sim P(\,\cdot\,|x_{i})$ is drawn independently, and $\mathbf{m}=(m_{1},\dots,m_{n})\in\mathcal{M}_{\rho}^{\,n}$ collects the masks for a mini-batch or dataset.
Given $\mathbf{m}$, define the block Gram matrices
$K_{U}(\mathbf{m})=\bigl{[}k_{u}(x_{i}^{U}(\mathbf{m}),\,x_{j}^{U}(\mathbf{m}))\bigr{]}_{i,j},$
$K_{M}(\mathbf{m})=\bigl{[}k_{m}(x_{i}^{M}(\mathbf{m}),\,x_{j}^{M}(\mathbf{m}))\bigr{]}_{i,j}.$
Let
$\Phi(\mathbf{m}):=\max_{\alpha,\beta}\frac{\alpha^{\!\top}K_{U}(\mathbf{m})\,K_{M}(\mathbf{m})\,\beta}{\sqrt{\alpha^{\!\top}K_{U}(\mathbf{m})^{2}\alpha}\;\sqrt{\beta^{\!\top}K_{M}(\mathbf{m})^{2}\beta}}$
be the optimal value of (7) for this mask realization. The population objective (6) is $\mathbb{E}_{\mathbf{m}}[\Phi(\mathbf{m})]$. Once $K_{U}(\mathbf{m})$ and $K_{M}(\mathbf{m})$ are known or estimated from a pilot sample, choosing a masking scheme amounts to selecting $P$ to put larger mass on masks with larger $\Phi(\mathbf{m})$, subject to the mask ratio and any structural constraints.
Linear scalar-kernel special case.
With linear kernels,
$k_{u}(x_{i}^{U},x_{j}^{U})=\Phi_{u}(x_{i}^{U})^{\top}\Phi_{u}(x_{j}^{U}),$
$k_{m}(x_{i}^{M},x_{j}^{M})=\Phi_{m}(x_{i}^{M})^{\top}\Phi_{m}(x_{j}^{M}),$
where $\Phi_{u}$ and $\Phi_{m}$ map views to feature vectors representing their semantics. It is recommended to choose $\Phi_{u},\Phi_{m}$ to be (approximately) alignment-invariant, so that views with similar semantics but different temporal indices (e.g., phase shifts/misalignment) still yield high kernel similarity.
In this case the sample cross-covariance and within-view covariance matrices are $\Sigma_{U\!U}:=\sum_{i=1}^{n}\Phi_{u}(x_{i}^{U})\,\Phi_{u}(x_{i}^{U})^{\top}/n$, $\Sigma_{M\!M}:=\sum_{i=1}^{n}\Phi_{m}(x_{i}^{M})\,\Phi_{m}(x_{i}^{M})^{\top}/n$ and $\Sigma_{U\!M}:=\sum_{i=1}^{n}\Phi_{u}(x_{i}^{U})\,\Phi_{m}(x_{i}^{M})^{\top}/n$. Then (7) reduces to the CCA problem (deep CCA if $\Phi$ is a deep network model):
$\max_{w_{u},w_{m}}\frac{w_{u}^{\top}\Sigma_{U\!M}w_{m}}{\sqrt{(w_{u}^{\top}\Sigma_{U\!U}w_{u})(w_{m}^{\top}\Sigma_{M\!M}w_{m})}},$ (9)
whose optimum is the top singular value $\sigma_{1}$ of $\Gamma=\Sigma_{U\!U}^{-1/2}\Sigma_{U\!M}\Sigma_{M\!M}^{-1/2}$ (Sánchez et al., 1979; Andrew et al., 2013). Thus, the optimal masking distribution concentrates on allocations yielding larger $\sigma_{1}(\Gamma)$. $\Gamma$ can be viewed as the biadjacency matrix of a weighted bipartite graph with unmasked and masked views as the two parts. Its top singular value $\sigma_{1}(\Gamma)$ is the spectral norm and quantifies the strongest global linear association across the two sides. A large $\sigma_{1}$ indicates strong cross-view structure (Sarkar & Dong, 2011).
4.6 Demonstration with Synthetic Data
To validate the above heuristic, we compare MoCA against the synchronized masking scheme through a simulation with synthetic data. Following the augmentation Algorithm 1, we generate synthetic data by sampling two samples $\bar{x}_{1},\bar{x}_{2}$ from the training set, and then randomly extract a consecutive chunk of time series from $\bar{x}_{2}$ to replace the corresponding time series in $\bar{x}_{1}$. The resulting synthetic time series represents a time window in which the participant transitions, e.g., from walking to sitting. We repeated this process 1000 times to obtain our simulation data. An example image is shown in Figure 1.
The MoCA and synchronized masking schemes yield different partitions of the images into unmasked and masked views. Under each masking scheme, we map the unmasked and masked views to feature vectors with an encoder then reduce to 50 dimensions with PCA (scores are centered by construction). Let $Z_{U}^{(s)},Z_{M}^{(s)}\in\mathbb{R}^{n\times 50}$ be the PC-score matrices. Define for each partition
$\Sigma_{U\!U}^{(s)}$ $=\tfrac{1}{n}(Z_{U}^{(s)})^{\top}Z_{U}^{(s)},$
$\Sigma_{M\!M}^{(s)}$ $=\tfrac{1}{n}(Z_{M}^{(s)})^{\top}Z_{M}^{(s)},$
$\Sigma_{U\!M}^{(s)}$ $=\tfrac{1}{n}(Z_{U}^{(s)})^{\top}Z_{M}^{(s)}.$
We then compute the matrix $\Gamma$ and calculate the top singular value $\sigma_{1}$. The resulting top singular values $\sigma_{1}^{\text{MoCA}}=0.925$ while $\sigma_{1}^{\text{Synchronize}}=0.836$.
The larger $\sigma_{1}$ under MoCA indicates a stronger cross-view association: MoCA produces partitions that better exploit cross-modal correlations and thus support improved reconstruction, while the slice-wise constraint of synchronized masking suppresses informative cross-modal pairings within a time window. Because a larger $\sigma_{1}$ implies a smaller lower bound for $L_{\mathrm{MAE}}$, MoCA is preferable to synchronized masking when within–time-slice correlations across modalities/axes are strong and should be leveraged, which is often the case in multi-modal wearable device data.